SIC Insurance Company bought a reinsurance product from a foreign reinsurance company in UK. The cost of the reinsurance product is £500,000 payable in 1 year time. Assume that the spot exchange rate is GHS5/£, and the 1-year forward rate is GHS5.5/£. The money market interest rate in Ghana is 15 percent and 10 percent in UK.

Required:

i. Describe how SIC can use a forward market hedge to manage this payable.

ii. Calculate the amount to undertake the forward market hedge?

iii. Calculate the gain or loss under the forward market hedge if the spot rate at maturity turns out to be GHS6/£.

iv. What option contract can SIC use to hedge this payable?

v. Suppose the spot rate at maturity is GHS4.7/£. Will SIC be better or worse off under the options market hedge? Assume that the premium is GHS0.001/£ and the strike price is GHS5.5/£.

vi. Describe how SIC can use a money market hedge to manage its exposure. Show the calculation of each step.
Answer #1

i & ii). Forward market hedge - Buy £500,000 forward 12 months at GHS5.5/£

Cost of settlement in 12 months will be 500,000*5.5 = GHS2,750,000

iii). If spot rate at maturity is GHS6/£ then GHS required would be 500,000*6 = GHS3,000,000

Amount saved (or gain) = 3,000,000 - 2,750,000 = GHS250,000

iv). It can be hedged by buying a call option on the £500,000 payable.

v). Since cost of capital for the company is not provided, we will use the money market 15% interest rate in its place.

Option premium = payable*0.001 = GHS500

Premium carried forward for a year = 500*(1+15%) = GHS575

If spot rate at maturity is GHS4.7/£ then the option won't be exercised, so GHS cost of payable will be

500,000*4.7 = GHS2,350,000

Net cost of call option hedge = 2,350,000 + 575 = GHS2,350,575

vi). Money market hedge: Exchange GHS for £ now and invest it for 12 months.

Payable required in 12 months = £500,000

Discount factor @ UK interest rate = (1+10%) = 1.10

Amount required now = 500,000/1.10 = £454,545.45

GHS at the current spot rate of GHS5/£ = 454.545.45*5 = GHS2,272,727.27

Amount after investing it at 15% for a year = 2,272,727.27*(1+15%) = GHS2,613,636.36 (Cost of settlement after a year)